OpenStudy (anonymous):

The measures of three angles of a quadrilateral are 80, 90, and 103 degrees. Find the measure of the fourth angle. 67 77 87 97

OpenStudy (anonymous):

@SpooksDF @Gralwyn

OpenStudy (anonymous):

the sum of the interior angles of a quadrilateral is 360 degrees

OpenStudy (anonymous):

so, 360 = 80 + 90 + 103 + x

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

The angles need to add up to 360 since a quadrilateral had 4 90 degree angles. So 80 + 90 + 103 + x = 360 273 + x = 360 Solve for X by subtracting 273 from both sides: x = 360 - 273

OpenStudy (anonymous):

So x = 87

OpenStudy (anonymous):

thnx:)